OpenStudy (anonymous):

Solve the system using substitution. 3x – y = -1 x – y = -3 What is the solution for y?

OpenStudy (amistre64):

if you want y, you need to replace x what is an equation for x using the second part?

OpenStudy (across):

You could also change the sign of the second equation and add it to the first.

OpenStudy (amistre64):

solution by method tho ... but yes, that would eliminate the ys and solve for x; and backtrack to get ys again

OpenStudy (anonymous):

x=y+3?

OpenStudy (amistre64):

very close, assuming a typo - x = y - 3

OpenStudy (amistre64):

lets use this equation in place of the "x" up top; that will give us just ys to play with 3x – y = -1 3(y-3) – y = -1

OpenStudy (amistre64):

the rest is just basic order of operations to get to y

OpenStudy (anonymous):

so then, it would be 3y-9-y=1 3y-8-y 4y=-8 y=-2

OpenStudy (amistre64):

lets see if we can follow that thru: 3y-9-y=1 <<< and this one needs to be = -1, careful with those 3y-8-y << this ones off a bit 3y-9-y=-1 : commute 3y-y-9=-1 : combine like terms 2y-9 = -1 : +9 both sides 2y = 8 : divide off the 2 y = 4

OpenStudy (anonymous):

oh

OpenStudy (amistre64):

keep up the practice ;)

OpenStudy (anonymous):

Thanks